A coolant widely used an automobiles engines is 38.70% carbon, 9.74% hydrogen, and 51.56% oxygen. It's molar mass is 62.07 g/mol . What is the molecular formula of the compound?
Answer #1

we have mass of each elements as:
C: 38.7 g
H: 9.74 g
O: 51.56 g


Divide by molar mass to get number of moles of each:
C: 38.7/12.01 = 3.2223
H: 9.74/1.008 = 9.6627
O: 51.56/16.0 = 3.2225


Divide by smallest to get simplest whole number ratio:
C: 3.2223/3.2223 = 1
H: 9.6627/3.2223 = 3
O: 3.2225/3.2223 = 1


So empirical formula is:CH3O

Molar mass of CH3O,
MM = 1*MM(C) + 3*MM(H) + 1*MM(O)
= 1*12.01 + 3*1.008 + 1*16.0
= 31.034 g/mol

Now we have:
Molar mass = 62.07 g/mol
Empirical formula mass = 31.034 g/mol
Multiplying factor = molar mass / empirical formula mass
= 62.07/31.034
= 2

So molecular formula is:C2H6O2
Answer: C2H6O2